Find the midpoint for A=(-6,6) B=(8,-8) don't understand how to apply formula and solve. Can someone give me the simpliest method for understanding.
Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

Find the midpoint for A=(-6,6) B=(8,-8)
----------
I don't know what formula you mean.
Find the average of x and y separately:
x: (-6 + 8)/2 = 1
y: (6 - 8)/2 = -1
--> (1,-1)

Answer by solver91311(24713)



The mid-point formulas:



and




where and

are the coordinates of the given points. That all means that to find the -coordinate of the mid-point, you take the -coordinate of one of your points and add it to the -coordinate of your other point, then divide the result by 2. Then you do the same thing with the -coordinates. Then you write the ordered pair to describe your mid-point.

John